A market researcher is studying the spending habits of people across age groups. The amount of money spent by each individual is classified by spending category (Dining out, Shopping, or Electronics) – Factor A, and generation (Gen-X, Gen-Y, Gen-Z, or Baby Boomers) – Factor B. An incomplete ANOVA table is shown below.

Source of Variation

SS

df

MS

F

Rows (Factor B)

Columns (Factor A)

46,869

Error

Total

1,321,831

35

How many degrees of freedom are there for factors A, B, and Error, respectively?

Group of answer choices

a. 3, 4, 28

b. 2, 3, 6

c. 2, 3, 30

d. 3, 4, 12
Answer #1

The given question ,solution is option {c},

I will explain

Factor B indicating rows so d.f is ( rows -1)=no of rows is (4-1=3)

And Factor A contain columns so d.f is (columns -1)= no of columns is (3-1=2)

And the final error = Total-c-r=35-3-2=30.

The final answer is (c)=2,3,30